OpenStudy (anonymous):

PLEASE HELP!!! MEDAL WILL BE REWARDED! Determine two pairs of polar coordinates for the point (2, -2) with 0° ≤ θ < 360° I got to ((radical225+pi^2)/3,0) but i cant figure out the rest of the problem!

OpenStudy (anonymous):

\[\sqrt{225 + \pi^2}/3, 0\]

OpenStudy (anonymous):



OpenStudy (anonymous):

you need \(r\) and \(\theta\)

OpenStudy (anonymous):

\(r=\sqrt{2^2+2^2}=\sqrt{8}=2\sqrt{2}\)

OpenStudy (anonymous):

Would it be (2 , (radical2), 225°), (-2 ,(radical2), 45°)

OpenStudy (anonymous):

now you have an infinite number of choices for \(\theta\)

OpenStudy (anonymous):

you can use \(\theta=-45\) or \(\theta=315\) or whatever

OpenStudy (anonymous):

or you can use what you said, \[(-2\sqrt{2},45)\]

OpenStudy (anonymous):

Okay, but that order would be correct? not the other way around? @satellite73

OpenStudy (anonymous):

yes, it is \((r, \theta)\)

OpenStudy (anonymous):

Okay thanks again :)

OpenStudy (anonymous):

yw